OpenStudy (anonymous):

Must show work! Find vertex, foci, directrix for (y-7)^2=-12(x+1)

OpenStudy (jdoe0001):

\(\large {\begin{array}{llll} (y&-{\color{red}{ k}})^2=&\ {\color{green}{ 4p}}(x&-{\color{blue}{ h}})\\ &\ \uparrow &\ \uparrow \\ (y&-{\color{red}{ 7}})^2=&-{\color{green}{ 12}}(x&{\color{blue}{ +1}}) \end{array}\\ \quad \\ vertex=(h,k)\\ p=\textit{distance form the vertex to the focus or directrix}}\)

OpenStudy (anonymous):

whats the answer though?

OpenStudy (jdoe0001):

well, what's the vertex?

OpenStudy (anonymous):

(1,7)

OpenStudy (anonymous):

(-1,7)

OpenStudy (jdoe0001):

sorry, I just happen to be a bit lagged =(

OpenStudy (jdoe0001):

so the vertex is at (-1, 7) and we know that -12 = 4p so what would be the "p" distance? well \(\bf -12=4p\implies -3=p\) so.... the negative means the graph is flipped the other way of the usual so.... notice we have a "squared y-coordinate" value that means the parabola is a horizontal one usually the positive one goes to the right so a negative number in front of the "p" means is opens or goes to the left 

OpenStudy (anonymous):

what are the foci?

OpenStudy (jdoe0001):

so the "p" is the distance form the vertex to the focus or directrix so p=3, so